# KAN4TSF: Are KAN and KAN-based models Effective for Time Series Forecasting?

### v1

@@ Abstract
Time series forecasting is a crucial task that predicts the future values of variables based on historical data. Time series forecasting techniques have been developing in parallel with the machine learning community, from early statistical learning methods to current deep learning methods. Although existing methods have made significant progress, they still suffer from two challenges. The mathematical theory of mainstream deep learning-based methods does not establish a clear relation between network sizes and fitting capabilities, and these methods often lack interpretability. To this end, we introduce the Kolmogorov-Arnold Network (KAN) into time series forecasting research, which has better mathematical properties and interpretability. First, we propose the Reversible Mixture of KAN experts (RMoK) model, which is a KAN-based model for time series forecasting. RMoK uses a mixture-of-experts structure to assign variables to KAN experts. Then, we compare performance, integration, and speed between RMoK and various baselines on real-world datasets, and the experimental results show that RMoK achieves the best performance in most cases. And we find the relationship between temporal feature weights and data periodicity through visualization, which roughly explains RMoK’s mechanism. Thus, we conclude that KAN and KAN-based models (RMoK) are effective in time series forecasting. Code is available at KAN4TSF: https://github.com/2448845600/KAN4TSF.
@@ 1 Introduction
Time series forecasting (TSF) is the task of using historical data to predict future states of variables. This research area includes a broad scope of applications, such as financial investment, weather forecasting, traffic estimation, and health management Bi et al. (2023); Gao et al. (2023); Savcisens et al. (2024); Han et al. (2024a). The machine learning community’s progress has long inspired time series forecasting technology: the popularity of early statistical learning methods gave rise to SVR and ARMIA, while the development of deep learning introduced MLP and Transformer into time series forecasting. At present, various time series forecasting methods cover almost all deep learning network architectures, such as RNN, CNN, Transformer, and MLP Nie et al. (2023); Wu et al. (2023); Han et al. (2024b). The forecasting models derived from different network architectures have their own advantages in forecasting performance, running speed, and resource usage.
Although deep learning-based models have made notable progress in time series forecasting, there are still several challenges. The universal approximation theorem (UAT), which is the mathematical foundation of most mainstream forecasting models, cannot provide a guarantee on the necessary network sizes (depths and widths) to approximate a predetermined continuous function with specific accuracy. And this theory can only achieve an approximation rather than a representation. The limitations of UAT have become the sword of Damocles hanging over time series forecasting. Furthermore, the prediction mechanism of existing models is black-box, resulting in a lack of interpretability. These nontransparent methods are suspected of being suitable for tasks that require a low tolerance for errors, such as medicine, law and finance.
Kolmogorov-Arnold Network (KAN) Liu et al. (2024a), which is based on the Kolmogorov-Arnold representation theorem (KART), has become a novel approach to solving the above challenges. On the one hand, KART proves that a multivariate continuous function can be represented as a combination of finite univariate continuous functions. This theorem establishes the relationship between network size and input shape under the premise of representation. On the other hand, KAN offers a pruning strategy that simplifies the trained KAN into a set of symbolic functions, enabling the analysis of specific modules’ mechanisms, thereby significantly enhancing the network’s interpretability. In addition, KAN’s function fitting idea is consistent with the properties of time series, such as periodicity and trend, which is conducive to embedding prior knowledge into the network structure and improving the performance of the network.
Despite being a relatively recent proposal, KAN, which employs a trainable 1D B-spline functions to convert incoming signals, has already sparked numerous efforts to improve or broaden its capabilities. Some studies propose KAN’s variants which replace the B-splines with Chebyshev polynomials SS (2024), wavelet functions Bozorgasl and Chen (2024), Jacobi polynomials Aghaei (2024), ReLU functions Qiu et al. (2024), etc., to accelerate training speed and improve network performance. Other studies introduce KAN with existing popular network structures for various applications. For example, ConvKAN Bodner et al. (2024) and GraphKAN Zhang and Zhang (2024); Xu et al. (2024) are proposed for image processing and graph processing. In summary, KANs have been extensively empirically studied in vision and language Azam and Akhtar (2024); Yu et al. (2024). However, existing studies lack a KAN-based model that considers time series domain knowledge, making it impossible to verify whether KAN is effective in time series forecasting.
To this end, we aim to propose a KAN-based model for the time series forecasting task and evaluate its effectiveness from four perspectives: performance, integration, speed, and interpretability. First, we propose the Reversible Mixture of KAN Experts model (RMoK), a KAN-based time series forecasting model that uses multiple KAN variants as experts and a gating network to adaptively assign variables to specific experts for prediction. RMoK is implemented as a single-layer network because we hope that it can have similar performance and better interpretability than existing methods. Then, we use a unified training and evaluation setting to compare the performance of RMoK and current popular baselines on seven real-world datasets. The experimental results show that RMoK achieves state-of-the-art (SOTA) performance in most cases. Subsequently, we conduct a comprehensive empirical study on KAN-based models, including the comparison between KAN and Linear, the effect of integrating KANs with the Transformer, and the speed of the KAN-based models. Finally, we discuss the interpretability of RMoK using the example of weather prediction. We visualize the weights of temporal features at different time steps in KAN and find the correlation between the weight distribution and the periodicity of the data.
To sum up, the contributions of this work include:
@@
  • To validate our claims, we propose the Reversible Mixture of KAN Experts model, which uses a single layer of the mixture of KAN experts to keep a balance between performance and interpretability.
  • We fairly compare the performance between RMoK and baselines on seven real-world datasets, and the experimental results show that RMoK achieves the best performance in most cases. And we also conduct a comprehensive empirical study on KAN-based models about integration and speed.
  • We mine the relationship between time feature weights and data periodicity through visualization, which roughly explains the mechanism of RMoK.
In summary, compared with the baselines in terms of performance, integration, speed, and interpretability, we conclude that KAN is effective in time series forecasting.
2 Problem Defintion
@@ 3 Related Work
3.1 Time-Series Forecasting Models
Figure 1: The computational process of Linear and KAN layers under a certain input and output dimension.
Although Transformer-based methods have almost become the standard in CV and NLP, various network architectures (such as Transformer, CNN, and MLP) are competing in time series forecasting recently.
Transformer-based time series forecasting models have strong performance but high time and memory complexity. Informer Zhou et al. (2021) proposed ProbSparse self-attention to reduce the complexity from $\mathcal{O}(T^{2})$ to $\mathcal{O}(T\log T)$. Pyraformer Liu et al. (2022a) utilizes the pyramid attention mechanism to capture hierarchical multi-scale time dependencies with a time and memory complexity of $\mathcal{O}(T)$. PatchTST Nie et al. (2023) and Crossformer Zhang and Yan (2023) use the patch operation to reduce the number of input tokens, thereby reducing time complexity. The performance of early Multi-layer perception-based (MLP-based) models is generally weaker than transformer-based methods. However, NLinear Zeng et al. (2023) and RLinear Li et al. (2023) combine different normalization methods with a single-layer MLP, and achieve performance that exceeds Transformer-based models on some datasets with extremely low computational cost.
Recurrent neural networks (RNNs) are suitable for handling sequence data, making them a favorable choice for time series analysis. The size of the hidden state in RNN is independent of the input time series length, so recent RNN-based time series prediction models, such as SegRNN Lin et al. (2023) and WITRAN Jia et al. (2024), apply the RNN structure to time series forecasting with longer input. Convolutional neural networks (CNNs) are frequently used in time series forecasting models in the form of 1D convolution, such as ModernTCN donghao and wang xue (2024) and SCINet Liu et al. (2022b), but TimesNet Wu et al. (2023) takes a different approach by converting the 1D time series into a 2D matrix through the Fourier transform and then using 2D convolution for prediction. In addition to the above four types of network structures, there is also time series forecasting work based on new network architectures such as Mamba Dao and Gu (2024); Wang et al. (2024) and RWKV Peng et al. (2023); Hou and Yu (2024).
@@ 3.2 Kolmogorov-Arnold Network
The Kolmogorov-Arnold representation theorem (KART) is the mathematical foundation of the Kolmogorov-Arnold Network (KAN) Liu et al. (2024a), which makes KAN more fitting and interpretable than Multi-Layer Perceptrons (MLP) based on the universal approximation theorem. We show the different between KAN and MLP in Figure 1. Given an input tensor $\mathbf{x}\in\mathbb{R}^{n_{0}}$, the structure of $L$ layers KAN network can be represented as:
@@
where $\phi$ consists two parts: spline function and the residual activation function with learnable parameters $w_{b},w_{s}$:
$\phi(x)=w_{b}\mathrm{SiLU}(x)+w_{s}\mathrm{Spline}(x),$ (3)
@@
Recent studies have proposed a large number of KAN variants to expand its application scope and improve its performance. Some work uses functions such as wavelet functions Bozorgasl and Chen (2024), Taylor polynomials SS (2024) and Jacobi polynomials Aghaei (2024) instead of spline functions to improve the performance of KAN in different fields. Some work extends KAN to the fields of vision and graphics and proposes ConvKAN and GraphKAN Zhang and Zhang (2024); Bodner et al. (2024). Few works try to apply KAN to time series, but there is a lack of sufficient theoretical and experimental analysis Vaca-Rubio et al. (2024). This paper aims to analyze the effectiveness of KAN and its variants in time series forecasting, we use XKAN to describe these methods uniformly (for example, TaylorKAN is KAN with Taylor polynomials).
4 RMoK
4.1 Mixture of KAN Experts Layer
Figure 2: The structure of RMoK and MoK layer.
Time series in real-world scenarios frequently exhibit non-stationarity, with their statistical properties (such as mean and variance) varying over time. Moreover, there are significant distribution discrepancies between variables in multivariate time series. This poses a significant challenge to time series forecasting techniques, inevitably impacting the KAN-based methods as well. Fortunately, when it comes to dealing with distribution shift, KAN has a unique characteristic compared to existing Linear and Transformer-based networks: KAN has many variants using different spline functions. Considering that the special spline function may be suitable for modeling certain data distributions, we try to combine several KANs into a single layer and adaptively schedule them according to the input data.
Following this idea, we propose the mixture of KAN experts (MoK) layer, which is a combination of KAN and mixture of experts (MoE). MoK layer uses a gating network to assign KAN layers to variables according to temporal features, where each expert is responsible for a specific part of the data. KAN and its variants only differ in the spline function in Equation 3, so we use $\mathcal{K}(\cdot)$ to represent these methods uniformly in this paper. Our proposed MoK layer with $N$ experts can be simple formed as:
@@ 4.1 Mixture of KAN Experts Layer
where $\mathcal{G}(\cdot)$ is a gating network. This mixture of experts structure can adapt to the diversity of time series, with each expert learning different parts of temporal features, thus improving performance on time series forecasting task.
The gating network is the key module of the MoK layer which is responsible for learning the weight of each expert from the input data. The softmax gating network $\mathcal{G}_{\mathrm{softmax}}$ which uses a softmax function and a learnable weight matrix $\mathbf{w}_{g}$ to schedule input data as:
@@
This gating network is popular due to its simple structure. However, it activates all experts once, resulting in low efficiency when there are a large number of experts. Therefore, we adopt the sparse gating network Shazeer et al. (2017) which only activates the best matching top-$k$ experts. It adds Gaussian noise to input time series by $\mathbf{w}_{\mathrm{noise}}$, and uses $\mathrm{KeepTopK}$ operation to retains experts with the highest $k$ values. The sparse gating network can be described as:
@@
Table 1: The statistics of the seven TSF datasets.

### v2

# Are KANs Effective for Multivariate Time Series Forecasting?

@@ Abstract
Multivariate time series forecasting is a crucial task that predicts the future states based on historical inputs. Related techniques have been developing in parallel with the machine learning community, from early statistical learning methods to current deep learning methods. Despite their significant advancements, existing methods continue to struggle with the challenge of inadequate interpretability. The rise of the Kolmogorov-Arnold Network (KAN) provides a new perspective to solve this challenge, but current work has not yet concluded whether KAN is effective in time series forecasting tasks. In this paper, we aim to evaluate the effectiveness of KANs in time-series forecasting from the perspectives of performance, integrability, efficiency, and interpretability. To this end, we propose the Multi-layer Mixture-of-KAN network (MMK), which achieves excellent performance while retaining KAN’s ability to be transformed into a combination of symbolic functions. The core module of MMK is the mixture-of-KAN layer, which uses a mixture-of-experts structure to assign variables to best-matched KAN experts. Then, we explore some useful experimental strategies to deal with the issues in the training stage. Finally, we compare MMK and various baselines on seven datasets. Extensive experimental and visualization results demonstrate that KANs are effective in multivariate time series forecasting. Code is available at: https://github.com/2448845600/EasyTSF.
@@ 1 Introduction
(a) Performance comparison
(b) Training Loss
Figure 1: Two experiental findings on the ETTh1 dataset. Subfigure (a) shows the univariate forecasting results of KANs with different base functions, where results are minmax-normalized and 0 means the minimum MSE in the row. Subfigure (b) reports the training loss of one/three layers of Linear/KAN models.
@@
The multivariate time series forecasting (MTSF) task aims to predict future states of variables from input historical data. Related technologies have been applied to a broad scope of fields, such as financial investment, weather forecasting, and traffic management Gao et al. (2023); Savcisens et al. (2024); Han et al. (2024). Time series forecasting methods have long been inspired by advances in the machine learning community. The popularity of early statistical learning methods gave rise to SVR and ARMIA Holt (2004); recurrent neural networks (RNN), which are specially designed to process sequence data, are naturally applied to time series analysis Hewamalage et al. (2021); TFT Lim et al. (2021) and Informer Zhou et al. (2021) promote the application of transformer architecture in time series prediction; LSTF-Linear Zeng et al. (2023) rethinks the value of Linear. At present, various deep learning network architectures, such as RNN, CNN, Transformer, and MLP, are widely used in time series forecasting Wang et al. (2023); Zhou et al. (2024); Jia et al. (2024). The forecasting models derived from different network architectures have their own advantages in forecasting performance, running speed, and resource usage.
However, most popular time series forecasting models are based on deep learning-based modules, whose prediction process is blackbox and prediction results lack interpretability. These nontransparent methods are suspected to be unsuitable for tasks with low error tolerance, such as medicine, law, and finance. Some traditional, non-deep learning-based methods are interpretable, but it is difficult to achieve the expected performance. Fortunately, a novel kind of deep learning network, the Kolmogorov-Arnold Network (KAN) Liu et al. (2024b), which is based on the Kolmogorov-Arnold representation theorem (KART), has become a feasible solution to the above dilemma of choosing between performance and interpretability. KAN offers a pruning strategy that simplifies the trained KAN model into a set of symbolic functions, making it possible to analyze the mechanisms of the network, thereby significantly enhancing the interpretability.
Although KAN has the advantages of both performance and interpretability in theory, some works Yu et al. (2024a) have pointed out that it does not work well for real-world tasks. A few works Genet and Inzirillo (2024); Vaca-Rubio et al. (2024) have tried to apply KAN to time series analysis, but they have not used the characteristics of time series data to design efficient KAN-based networks, and lack fair comparisons on common datasets. Existing works cannot provide a convincing conclusion on whether KAN is effective in time series forecasting. Therefore, we aim to provide a reliable benchmark to comprehensively evaluate the effectiveness of the KAN in multivariate time series forecasting tasks for the first time. When applying KAN to time series forecasting, we encountered two challenges, as following.
One is how to select the most suitable KAN. The KAN proposed by Ziming Liu et al. is not only a model but also a general function-fitting framework. Subsequent studies SS (2024); Bozorgasl and Chen (2024); Aghaei (2024) use different base functions to replace the B-spline function for better performance on specific tasks. We report the experimental results about five KAN’s variants (KANs), each with a special base function, on seven univariates of the ETTh1 dataset, as illustrated in Figure 1(a). This makes it cumbersome to determine the appropriate KAN for achieving the best performance. The other is how to train the deeper KAN. We build a single-layer KAN model that can achieve similar performance with a single-layer Linear model. But as the layer of the KAN model increases, the performance drops significantly. We visualize the values of loss in 10 times repeated trainings in Figure 1(b) and find that the initial loss of KAN is larger than that of Linear, and the loss of the 3-layer KAN model becomes significantly larger and more unstable.
To address the above challenges, we first apply the mixture-of-experts technique to KAN and obtain the Mixture-of-KAN (MoK) layer, which uses a gating network to adaptively assign the input to the matching KAN layers with special base functions according to the variable dimensions. This structure aggregates the advantages of KAN’s variants in different cases. Then, we build a simple yet efficient Multi-layer Mixture-of-KAN network (MMK) based on MoK. By using modules that do not compromise the ability to be converted into symbolic functions, MMK achieves excellent performance while retaining the interpretability from KAN. Next, we analyze the phenomenon of abnormal training loss and use some training strategies to improve the performance of the proposed MMK. Finally, we evaluate the effectiveness of KANs in multivariate time series forecasting from four perspectives: performance, integration, speed, and interpretability. We compare MMK with many popular time series forecasting models on seven commonly used datasets, and experimental results show that MMK achieves state-of-the-art performance in most cases. We integrate MoK into recently proposed models and find that it can improve their performance. And we visualize the output weights of the gating network in the MoK layer and analyze the connection between KAN’s symbolic function conversion ability and domain knowledge.
To sum up, the contributions of this paper include:
@@
  • To validate our claims, we propose an interpretable Multi-layer Mixture-of-KAN network (MMK), which uses the mixture-of-KAN layer to aggregate the advantages of KAN’s variants. And we use some training strategies to improve performance.
  • We fairly compare MMK with many baselines on seven real-world datasets, and the experimental results show that MMK achieves the competitive performance.
In summary, we can conclude that KANs are effective in multivariate time series forecasting.
@@ 2 Related Work
2.1 Multivariate Time-Series Forecasting
Various network architectures (such as Transformer, CNN, and MLP) are competing in time series forecasting recently. Transformer-based time series forecasting models have strong performance but high time and memory complexity. Informer Zhou et al. (2021) proposed ProbSparse self-attention to reduce the complexity from $\mathcal{O}(T^{2})$ to $\mathcal{O}(T\log T)$. Pyraformer Liu et al. (2022b) utilizes the pyramid attention mechanism to capture hierarchical multi-scale time dependencies with a time and memory complexity of $\mathcal{O}(T)$. PatchTST Nie et al. (2023) and Crossformer Zhang and Yan (2023) use the patch operation to reduce the number of input tokens, thereby reducing time complexity. The performance of early Multi-layer perception-based (MLP-based) models is generally weaker than transformer-based methods. However, NLinear Zeng et al. (2023) and RLinear Li et al. (2023) combine different normalization methods with a single-layer MLP, and achieve performance that exceeds Transformer-based models on some datasets with extremely low computational cost. Recurrent neural networks (RNNs) are suitable for handling sequence data, making them a favorable choice for time series analysis. The size of the hidden state in RNN is independent of the input time series length, so recent RNN-based time series prediction models, such as SegRNN Lin et al. (2023) and WITRAN Jia et al. (2024), apply the RNN structure to time series forecasting with longer input. Convolutional neural networks (CNNs) are frequently used in time series forecasting models in the form of 1D convolution, such as ModernTCN donghao and wang xue (2024) and SCINet Liu et al. (2022a), but TimesNet Wu et al. (2023) takes a different approach by converting the 1D time series into a 2D matrix through the Fourier transform and then using 2D convolution for prediction.
@@ 2.2 Kolmogorov-Arnold Network
The Kolmogorov-Arnold representation theorem (KART) is the mathematical foundation of the Kolmogorov-Arnold Network (KAN) Liu et al. (2024b), which makes KAN more fitting and interpretable than Multi-Layer Perceptrons (MLP) based on the universal approximation theorem. Given an input tensor $\mathbf{x}\in\mathbb{R}^{n_{0}}$, the structure of $L$ layers in a KAN network can be represented as:
@@
where $\phi$ consists two parts: the spline function and the residual activation function with learnable parameters $w_{b},w_{s}$:
$\phi(x)=w_{a}\mathrm{SiLU}(x)+w_{b}\mathrm{Spline}(x),$ (3)
@@
Recent studies have proposed many variants of KAN to expand its application scope and improve its performance. Some work uses functions such as wavelet functions Bozorgasl and Chen (2024), Taylor polynomials Yu et al. (2024b) and Jacobi polynomials Aghaei (2024) instead of spline functions to improve the performance of KAN in different fields. We can unify these KANs using the generalized version of Equation 3 as:
$\phi(x)=w_{a}\mathbb{A}(x)+w_{b}\mathbb{F}(x).$ (4)
3 Multi-layer Mixture-of-KAN Network
3.1 Problem Defintion
3.2 Mixture-of-KAN Layer
Real-world time series data frequently exhibit non-stationarity, with their statistical properties (such as mean and variance) varying over time. Moreover, there are significant distribution discrepancies between variables in multivariate time series. This poses a significant challenge to time series forecasting techniques, inevitably impacting the KAN-based methods as well. Fortunately, when it comes to dealing with distribution shift, KAN has a unique characteristic compared to existing methods: it has many variants using different base functions. Considering that the special base function may be suitable for modeling certain data distributions, we try to combine several KANs into a single layer and adaptively schedule them according to the input data.
Following this idea, we propose the Mixture-of-KAN (MoK) layer, which is a combination of KAN and mixture-of-experts (MoE). The MoK layer uses a gating network to assign KAN layers to variables according to temporal features, where each expert is responsible for a specific part of the data. KAN and its variants only differ in the base function, so we use $\mathcal{K}(\cdot)$ to represent KANs uniformly in this paper. Our proposed MoK layer with $N$ experts can be simply formed as:
@@ 3.2 Mixture-of-KAN Layer
where $\mathcal{G}(\cdot)$ is a gating network. This mixture of expert structure can adapt to the diversity of time series, with each expert learning different parts of temporal features, thus improving performance on time series forecasting task.
The gating network is the key module of the MoK layer, which is responsible for learning the weight of each expert from the input data. The softmax gating network $\mathcal{G}_{\mathrm{softmax}}$ which uses a softmax function and a learnable weight matrix $\mathbf{w}_{g}$ to schedule input data as:
@@
This gating network is popular due to its simple structure. However, it activates all experts once, resulting in low efficiency when there are a large number of experts. Therefore, we adopt the sparse gating network Shazeer et al. (2017) which only activates the best matching top-$k$ experts. It adds Gaussian noise to input time series by $\mathbf{w}_{\mathrm{noise}}$, and uses $\mathrm{KeepTopK}$ operation to retain experts with the highest $k$ values. The sparse gating network can be described as:
@@
Figure 2: The architecture of MMK. Subfigure (a) is the structure of the mixture-of-KAN (MoK) layer; subfigure (b) is the pre-sample initialization strategy for the last KAN layer.
3.3 Network Architecture
The interpretability of KAN comes from the fact that the trained KAN model can be converted into a combination of many simple symbolic functions. In order to maintain this advantage, we use layers or operators with the same properties to build our KAN-based forecasting model.
RevIN
RevIN Kim et al. (2022) includes symmetrically structured normalization $\text{RevIN}^{+}$ and denormalization operations $\text{RevIN}^{-}$ in the variable dimension. Before sending data $X$ to the backbone network, $\text{RevIN}^{+}$ uses a learnable affine transformation to normalize the input time series of each variable. After the last MoK layer, outputs are denormalized to the original distribution space using the same affine transformation parameters in the first step by $\text{RevIN}^{-}$ operation. Each learnable parameter of RevIN corresponds to the statistical value of the time series variable, and it can be viewed as a symbolic function in form, so it does not harm the interpretability of the overall network.
MoK Block
Inspired by the ResNet He et al. (2016), we design a simple MoKBlock, which consists of a MoK layer, BatchNorm, and Dropout in sequence, and adds a residual structure. By stacking several MoKBlocks, MMK can obtain the stronger feature extraction capability.
$\operatorname{MoKBlock}(\boldsymbol{h})=\operatorname{Dropout}(\operatorname{BN}(\boldsymbol{h}+\operatorname{MoK}(\boldsymbol{h}))).$ (9)
Loss Function
During the training stage, the gating network has a tendency to reach a winner-take-all state where it always gives large weights for the same few experts. Following the previous work Shazeer et al. (2017), we add a load balancing loss function to encourage experts have equal importance. First, we count the weight of experts as loads, and calculate the square of the coefficient of variation ($\operatorname{CV}(\cdot)$) of the load values as additional loss:
$\mathcal{L}_{\text{load-balancing}}=\operatorname{CV}(\text{loads})^{2}.$ (10)
With this loss function, all experts are encouraged to have equal importance in the training stage. The total loss function is the sum of the prediction loss and load balancing loss:
$\mathcal{L}=\mathcal{L}_{\text{pred}}+\mathcal{L}_{\text{load-balancing}},$ (11)
where $\text{where }\mathcal{L}_{\text{pred}}=\operatorname{MSE}(\mathcal{Y},\hat{\mathcal{Y}}).$
3.4 Training Strategy
Although our proposed MMK is carefully designed, the performance of multiple-layer MMK is not as good as that of a single-layer MMK. From the visualization of the training loss, we infer that the main issue in training deep MMK is that the initial loss is too large and unstable. In the original KAN paper, $w_{a}$ is initialized according to the Xavier initialization and $w_{b}$ is initialized to 1 to guarantee variance-preserving Glorot and Bengio (2010); He et al. (2015); Yang and Wang (2024) with the assumption that all $x_{i}$ in $x$ are mutually independent. However, this initialization method is unstable when the number of network layers increases, because its premise, the mathematical expectation remains 0, is heavily dependent on the distribution of input data. We compare expectations of KAN and Linear in Equation 12:
$\mathbb{E}(\boldsymbol{y})^{\text{KAN}}$ $=\mathbb{E}[\phi(\boldsymbol{x})]$ (12)
$=\mathbb{E}[\boldsymbol{w}_{a}]\mathbb{E}[\mathbb{A}(\boldsymbol{x})]+\mathbb{E}[\boldsymbol{w}_{b}]\mathbb{E}[\mathbb{F}(\boldsymbol{x})],$
$=\mathbb{E}[\mathbb{F}(\boldsymbol{x})],$
$\mathbb{E}(\boldsymbol{y})^{\text{Linear}}$ $=\mathbb{E}[\boldsymbol{w}\boldsymbol{x}]=\mathbb{E}[\boldsymbol{w}]\mathbb{E}[\boldsymbol{x}]=0.$
The Linear layer keeps $\mathbb{E}=0$ by initializing $\boldsymbol{w}=0$, while KAN keeps $\mathbb{E}=0$ by supposing $\mathbb{E}[\mathbb{F}(\boldsymbol{x})]=0$. Since the real-world time series data has a biased distribution and non-negligible number of outliers, it is difficult to maintain $\mathbb{E}[\mathbb{F}(\boldsymbol{x})]=0$. Therefore, we propose a pre-sampling initialization strategy: before training starts, we forward all training samples to the model, and record the $\operatorname{Var}[\boldsymbol{x}]$ and $\operatorname{Var}[\mathbb{F}(\boldsymbol{x})]$ to initialize the $\boldsymbol{w}_{b}$ with $\mathcal{N}(0,\frac{\operatorname{Var}[\mathbb{F}(\boldsymbol{x})]}{\operatorname{Var}[\boldsymbol{x}]})$.